OpenStudy (greatlife44):

Suppose that the cost function for a certain product is 𝐶(𝑥) = 400 + 2𝑥 and the price function is 𝑝(𝑥) = 20 − 0.02𝑥. Find the production level that maximizes profit.

OpenStudy (studygurl14):

First you need to find profit. Subtract the cost from the price to get profit

OpenStudy (greatlife44):

So essentially, \[Profit = Price - Cost \] \[f(x) = p(x)-c(x)\]

OpenStudy (studygurl14):

yes

OpenStudy (studygurl14):

This is of course, assuming cost means cost to the producer, and price means how much money it takes for someone to buy it

OpenStudy (greatlife44):

20-0.02x - (400+2x) 20-0.02x -400 - 2x Combining like terms 20-(-400) 0.02x-2x I get this equation for profit. y(x) = -2.02x+420

OpenStudy (studygurl14):

it would be -380, not +420

OpenStudy (greatlife44):

I know what to do after, you find the derivative and set = to zero

OpenStudy (greatlife44):

y(x) = -380-2.02x

OpenStudy (studygurl14):

yes

OpenStudy (greatlife44):

\[\frac{ dy }{ dx }(-380-2.02x) = -2.02 \]

OpenStudy (greatlife44):

To me this answer doesn't make much sense.

OpenStudy (greatlife44):

@hartnn

OpenStudy (anonymous):

this doesn't have a largest value is there supposed to be an \(x^2\) somewhere in the question?

OpenStudy (greatlife44):

:( well... it didn't say this in the question but I found in my notes that revenue R(x) = x*p(x) or revenue = x times the price function

OpenStudy (dumbcow):

yes .... profit = revenue - cost = x*price - cost

OpenStudy (anonymous):

that makes more sense

OpenStudy (greatlife44):

Okay this is what I got now \[x*p(x)-C(x) = x*(20-0.02x)-400+2x \] \[20x-0.02x^{2}-400+2x\]

OpenStudy (anonymous):

\[x( 20 − 0.02𝑥)-( 400 + 2𝑥)\] looks better

OpenStudy (anonymous):

no, don't forget to distribute the minus sign!

OpenStudy (greatlife44):

oh

OpenStudy (greatlife44):

Okay so would it be this? \[20x-0.02x^{2}-400-2x = -0.02x^{2}+18x-400\]

OpenStudy (greatlife44):

wondering if I grouped the terms properly

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

max is at \(-\frac{b}{2a}=-\frac{18}{2\times (-.02)}\)

OpenStudy (greatlife44):

okay, is that a general formula for any parabolas -b/2a for the maximum?

OpenStudy (greatlife44):

the way I was taught was you take the derivative and set it = 0

OpenStudy (anonymous):

max or min, depending if it opens up or opens down it is the first coordinate of the vertex

OpenStudy (greatlife44):

so if it's value is >0 it opens up, maximum <0 down minimum?

OpenStudy (anonymous):

on the other hand, the derivative of \[y=ax^2+bx+c\] is \[y'=2ax+b\] set it equal to zero and solve, you always get \[x=-\frac{b}{2a}\]

OpenStudy (anonymous):

and no, you have that backwards, if it opens down you have a max, if it open up, you have a min

OpenStudy (anonymous):



OpenStudy (greatlife44):

interesting -0.04x+18 = 0 -18/-0.04 x = 18/0.04 x = 450 so this is what I got so I need to put this value back into original equation to find out what the maximum profit.

OpenStudy (greatlife44):

Oh I see now, the figure you drew clarified it.

OpenStudy (anonymous):

good

OpenStudy (greatlife44):

So now that i got the derivative I take what I get and plug it into the original function we started out with, and this gives us the profit. -0.02(450)^{2}+18(450)-400 = $3,650 So I also thought to plug in what I got for the other equations too p(x) = 20-0.02(450) = 11 so the price of each unit is 11 dollars and the cost C(x) = 400+2(450) = $1,300 my prof wrote down to test if its a local maximum or minimum but what's the point of that anyway cant we just find the derivate at zero and plug it back into f(x)?

OpenStudy (anonymous):

since you know it is a parabola that opens down, you know it has a max and not a min not sure what else there is to say

OpenStudy (greatlife44):

ok

OpenStudy (greatlife44):

so the profit is 3,650 the unit price is 11 dollars and it will cost 1,300